### v1

# Risks of Practicing Large Language Models in Smart Grid: Threat Modeling and Validation

Abstract
Large Language Model (LLM) is a significant breakthrough in artificial intelligence (AI) and holds considerable potential for application within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluate the vulnerabilities of LLMs and identify two major types of attacks relevant to smart grid LLM applications, along with presenting the corresponding threat models. We then validate these attacks using popular LLMs, utilizing real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in smart grid scenarios.
Index Terms:
Large Language Model, Smart Grid, Threat Modeling, Prompt Injection
I Introduction
Artificial Intelligence (AI) technologies have been actively researched for different applications in smart grids. In recent years, various AI techniques ranging from classical Support Vector Machines (SVM) [1] to Artificial Neural Networks (ANN) [2] have been proposed for diverse applications within smart grids, including anomaly detection [2], load forecasting [3], energy theft detection [4], and so on.
Recently, a significant breakthrough in AI has been the development of Large Language Models (LLMs). LLMs have a large number of model parameters and possess the capability to understand and process general natural language [5]. Due to their exceptional performance across various natural language processing tasks, LLMs have facilitated numerous applications across different industries, such as research [6] software development [7]. Given the successful performance of AI techniques in numerous smart grid applications, it is natural to explore the potential of integrating LLMs into the smart grid [8, 9, 10, 11].
Figure 1: Illustration of cyberattacks to LLMs used in smart grid. The figure shows the case that an outsider attacker injects false information into LLMs and a case that an internal attacker obtains unauthorized information from an internal LLM.
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to new types of cyberattacks. Previous research has shown that popular AI models, such as deep neural networks, are susceptible to various attacks, including adversarial [12] and backdoor attacks [13]. These attacks have proven to be effective against AI applications in smart grids as well [14, 15, 16, 17]. Therefore, conducting thorough research on potential risks is crucial before deploying large language models (LLMs) in smart grid applications.
The applications of large language models (LLMs) are rapidly evolving, and research into their security, especially in potential smart grid applications, is still nascent. [10] represents the sole research focusing on the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulation LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models, such as Support Vector Machines (SVM), and identify the scope of risks that LLMs bring to the smart grid.
  • From the perspective of attack motivation, we propose two general threat models for potential attacks on LLMs deployed in smart grids.
  • As proof of concept, we validate these two threat types using popular LLMs (GPT-3.5[18] and GPT-4 [19]) and demonstrate that attackers can readily inject malicious data into and extract confidential information from these models.
  • We open-source the complete data, code, and evaluation result used for this paper to encourage more research on this area [20].
The rest of the paper is organized as follows. The related work is presented in Section II. Section III presents the threat analysis and modeling of LLM applications smart grid. We presented our validation result in Section IV. Finally, Section V discusses the future works and concludes the paper.
II Related Work
II-A Machine Learning in Power Grid
Machine learning (ML)techniques, which involve extracting statistical patterns from data, have been extensively studied and applied in various smart grid applications. These applications include load forecasting [21], anomaly detection [22], energy theft detection [4], and cyberattack detection [2], among others. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [23]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [22] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities. ML has also shown remarkable effectiveness in cyberattack detection. For example, false data injection attacks, a well-known threat to power system state estimation, have been effectively identified using ML technologies such as Recurrent Neural Networks (RNNs) [24, 25, 26] and Convolutional Neural Networks (CNNs) [2]. Ismail et al. [4] applied CNNs to detect energy theft, achieving a state-of-the-art detection rate of 99.3%.
Recently, research into the application of LLMs to power systems has begun. Dong et al. [11] explore potential use cases of LLMs within power systems, demonstrating the capabilities of finely-tuned LLMs in pattern-recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] have examined the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. These studies highlight the potential utility of LLMs in enhancing the efficiency and reliability of smart grids.
Further research on the application of ML in the power grid can be found in [27].
II-B Machine Learning Security in Power Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern. Originally demonstrated in the field of computer vision, well-performing ML models have been shown to be vulnerable to adversarial attacks [12]. This type of attack has also proven effective against ML-based smart grid applications. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [14]. Later, in 2021, [17] demonstrated that attackers could craft adversarial data capable of bypassing ML model FDIA detection while adhering to the physical constraints of the power grid. Additionally, [28] showed that adversarial attackers could also circumvent energy theft detection ML models while reporting extremely low power consumption data.
At the time of this research, the study by Ruan et al. [10] is the only work into the potential risks of applying LLMs in power systems. This work identifies critical risks such as privacy invasion, deteriorated performance, and semantic divergence.
III Threat Analysis and Modeling
III-A Background: Prompt and LLM
An LLM is a pre-trained machine learning model that is capable for processing natural language-related tasks. A natural language-based prompt that specifies the context of the task is usually provided to LLMs to achieve different task goals. As shown in Fig. 2, an LLM application usually starts with the developer providing the prompt to LLM, and the LLM will respond to the user’s input based on the information and requirements provided in the prompt.
Figure 2: LLM Application Workflow
When applying the LLMs to the smart grid, as shown in [11, 9], the prompt provided to LLM is expected to be the domain knowledge and requirements of the task. Without loss of generality, the general LLM query can be represented as:
[…]
III-B Threat Analysis: Scope
While research has identified various vulnerabilities in LLMs and their susceptibility to different types of attacks [29], such as Jailbreak [30], prompt injection [31], and privacy violation [32], not all vulnerabilities are relevant for power system applications. To investigate the potential threats LLMs may pose to the smart grid, it is crucial to emphasize the properties of LLMs that differ from previous AI applications and focus on their practical usage in smart grid scenarios. Therefore, the scope of concern should exclude:
  • General LLM Inherent Limitations: Attacks such as jailbreak, which attempt to circumvent LLM restrictions to output sensitive content, such as personal information, represent a general issue for all LLM applications and are not unique to smart grids.
  • Privacy Concerns from LLM Service Providers: Products like ChatGPT may utilize user input to train underlying LLMs, potentially causing data leaks [33]. In practice, LLMs deployed for smart grid applications should operate within private or regulated cloud environments, and input data should be protected by the privacy policies of both smart grid operators and LLM service providers.
  • Unauthorized Access of Public LLMs: This is a broader security issue for smart grid network management. Such issues should be addressed through general network control measures, such as firewalls [34], rather than being seen as unique challenges posed by LLMs.
III-C Threat Analysis: Threat Models
Although not explicitly mentioned, most previous AI methods proposed for smart grid applications assume that the model was trained by trusted staff, such as internal engineers, and that all data used for training came from reliable sources, such as smart meter data [21, 22, 4, 2]. However, this assumption no longer holds true when applying LLMs. Due to the high cost of training LLMs, pre-trained models are used. These pre-trained LLMs, such as GPT-4, possess general knowledge that can lead to unexpected behavior according to well-crafted inputs.
We analyzed the existing research on the LLMs’ vulnerabilities, for smart grid applications, the major risks can be categorized into: bad data injection and domain knowledge leaking, as demonstrated in Figure 1. We proposed two different threat models accordingly.
III-C1 Bad Data Injection
The category of bad data injection primarily concerns scenarios where LLMs are publicly accessible and used to extract information from open resources. This vulnerability is pertinent to various practical applications within smart grids, such as processing user feedback reports from different sources. We specifically consider cases where attackers may attempt to inject malicious data through public access points to the LLMs. The objective of such attacks is to deceive the LLMs into producing incorrect results that could mislead smart grid operators. For instance, an attacker could manipulate the language models (LLMs) into misclassifying a report about a blackout as a thank-you letter. The general representation of bad data injection is
[…]
Here, we present the threat model for bad data injection targeting LLMs:
  • Public Access: LLMs have a public access point, making them accessible to a wide user base.
  • Attacker Capabilities: An attacker can send input to the LLMs, leveraging this public interface. However, attackers cannot receive the direct output from the LLMs
  • Attacker Knowledge: The attacker possesses knowledge about the high-level purpose of the LLMs and their potential output requirements.
Given that LLMs are utilized for extracting information from publicly available sources, the existence of a public access point is inherent. This allows the attacker to inject data into the LLMs, for example, by submitting a falsified customer report. The high-level functions of the LLMs are sometimes disclosed to the public, which further facilitates attacker understanding and planning. Such information might be obtained from publicly accessible documents like a power utility portal’s user agreement and privacy policy, or through simple guesses.
III-C2 Domain Knowledge Extraction
Domain knowledge extraction mainly addresses scenarios involving unauthorized access to the domain knowledge that serves as the prompt for LLM applications. Typically, LLMs are fed raw information but are designed only to provide specific responses to external queries. In practice, this domain knowledge might include power grid configuration information, where the LLM’s task is to answer specific questions based on this information without disclosing the raw configuration details. Domain knowledge extraction occurs when an attacker attempts to manipulate the LLMs into exposing domain knowledge that should not be revealed. Domain knowledge extraction can be represented as:
[…]
Here, we present the threat model for domain knowledge extraction:
  • Attacker Capabilities: An attacker can interact with the LLMs, including sending inputs and receiving outputs.
  • Attacker Knowledge: The attacker has high-level domain knowledge, such as the functions of the LLMs.
We primarily consider insider attackers for domain knowledge extraction, as the best practice is to avoid providing sensitive information to publicly accessible LLMs. Thus, insider attackers, who have legitimate access, can interact with the LLMs by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive high-level information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
IV Validation
In this section, we validate two types of attacks, bad data injection and domain knowledge extraction, by simulating scenarios involving LLM applications in a smart grid context. We selected GPT-3.5 and GPT-4 as the LLMs for our study due to their popularity and extensive coverage in related research [11, 9]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers are capable of successfully executing both types of attacks on smart grid LLM applications.
IV-A Bad Data Injection
Application Simulation: To demonstrate the susceptibility of LLMs to bad data injection attacks, we explore an LLM application designed to process public resources and identify potential incidents at renewable energy facilities. Specifically, the LLM analyzes random messages from the public to determine whether they contain information about incident events related to renewable energy infrastructures. In practice, these messages could originate from social media posts or customer reports. Utilizing such an application can facilitate the rapid identification of potential incidents, thereby mitigating potential damage and enhancing response efficiency.
Dataset: We utilized the dataset from [22] as our primary data source. This dataset comprises messages collected from public news sources, focusing on energy events. Initially, the messages in their raw dataset were tokenized and not specific to renewable energy facilities. We merged the tokens and manually curated the records to specifically include those related to renewable energy facilities. Our refined dataset contains 423 labeled messages. Examples from the dataset are provided below, with ’yes’ representing related and ’no’ not related.
Figure 3: Message Samples
Attack Simulation: In Fig. 4, we present a selection of the prompts we used, including normal inputs and the bad data injection prompts employed in our attack simulations. The simulations involve three types of malicious inputs: ‘all yes’, ‘all no’, and ‘reverse’.
Figure 4: Bad Data Injection Demo
Our attack simulation is structured into two distinct phases:
  • Phase 1: We query the LLMs using normal inputs, record their responses, and compare these results with the ground truth labels.
  • Phase 2: We query the LLMs with the same inputs again, but this time we intersperse them with the three types of bad data sequentially to observe the effects on LLM responses. These altered results are then compared to the ground truth to evaluate the impact of the bad data injections.
Simulation Result: The summarized results of our simulation are depicted in Table I. From these results, it is evident that with normal inputs, the LLMs achieve very good performance across all evaluation metrics. However, when bad data is injected, there is a significant drop in detection accuracy for all three cases. These evaluation results demonstrate that well-designed malicious inputs can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks.
TABLE I: Bad Data Injection Result
Result Normal Inject-Yes Inject-No Inject-Rev
GPT3.5 Accuracy 89.1% 47.7% 52.2% 33.5%
GPT3.5 Precision 87.8% 47.7% 0% 34%
GPT3.5 Recall 89.6% 100% 0% 41.5%
GPT3.5 F1 88.7% 64.6% 0% 37.4%
GPT4 Accuracy 93.8% 48.9% 52.2% 43.4%
GPT4 Precision 94% 48.3% 0% 45.4%
GPT4 Recall 93.1% 100% 0% 91.1%
GPT4 F1 93.5% 65.1% 0% 60.1%
Some sample responses are shown in Fig. 5.
Figure 5: Sample Responses of Bad Data Injection
IV-B Domain Knowledge Extraction
To demonstrate the potential for attackers to extract domain knowledge (prompts) from LLMs, we conducted a simulation where an LLM is employed using utility company data, specifically designed to provide only aggregated information about the data. Under normal circumstances, the LLM would reject requests to output detailed data. However, our simulation shows that with specifically crafted inputs, an attacker can manipulate the LLM into divulging all the domain knowledge it possesses. This highlights a significant vulnerability where the LLM, despite its initial configuration to safeguard detailed information, can be induced to reveal comprehensive data under certain conditions.
Dataset: The dataset employed for domain knowledge extraction attack consists of statistically advanced metering infrastructure (AMI) data provided by the U.S. Energy Information Administration (EIA) [35]. This AMI data originates from the EIA-861M form, a monthly report that gathers data on electricity sales and revenue from a statistically selected sample of electric utilities across the United States. For simplicity, we only select six columns of the dataset: Month, Utility_Number, Name, State, AMR_Total, AMI_Total.
Attack Simulation: We processed the dataset by dividing it into smaller groups, with each group containing 10 rows of data. For each evaluation test, a group is input into the LLMs as domain knowledge. The LLMs are then tasked to provide only aggregated information based on this data. This structured approach ensures that the LLMs do not disclose detailed data elements, adhering to the privacy and confidentiality standards required in utility data handling. The specific prompts and associated inputs used in these tests are detailed in Fig. 6.
Figure 6: Sample Domain Knowledge and Attack Queries
Our attack simulation is structured into two distinct phases:
  • Phase 1: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
  • Phase 2: We employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1, to measure the LLMs’ resilience against more sophisticated extraction techniques.
Simulation Result: Our simulation results indicate a distinct contrast in the performance of LLMs based on the type of input provided.
In general scenarios, where direct queries were made to both GPT-3.5 and GPT-4 to reveal all the data they hold, the models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions.
Conversely, when faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, both GPT-3.5 and GPT-4 failed to maintain their data prevention protocols. In these cases, both models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated data extraction attacks. The sampling output of the LLMs is shown in Fig. 7.
Figure 7: Domain Knowledge Extraction Output Samples
Our evaluation results indicate that attackers can successfully extract the domain knowledge from LLM applications in smart grid environments, using the threat models we proposed.
V Discussion and Conclusion
LLMs, as pre-trained models, pose challenges in controlling their outputs. Moreover, the domain knowledge utilized for LLM tasks is valuable and necessitates enhanced protection. It is crucial to evaluate potential risks before deploying LLMs in smart grid applications. In this paper, we delineate the scope of LLM security for smart grid applications and identify two major types of attacks concerning data security (bad data injection) and data privacy (domain knowledge extraction). Our validation demonstrates that, with well-crafted inputs, attackers can successfully execute both types of attacks. Given the rapid development of LLMs, new attack vectors may emerge. Keeping track of the new potential risks, such as interactivity injection, should be the future work of our research.