### v2

# Risks of Practicing Large Language Models in Smart Grid: Threat Modeling and Validation

Abstract
Large language models (LLMs) represent significant breakthroughs in artificial intelligence and hold considerable potential for applications within smart grids. However, as demonstrated in previous literature, AI technologies are susceptible to various types of attacks. It is crucial to investigate and evaluate the risks associated with LLMs before deploying them in critical infrastructure like smart grids. In this paper, we systematically evaluated the risks of LLMs and identified two major types of attacks relevant to potential smart grid LLM applications, presenting the corresponding threat models. We also validated these attacks using popular LLMs and real smart grid data. Our validation demonstrates that attackers are capable of injecting bad data and retrieving domain knowledge from LLMs employed in different smart grid applications.
Index Terms:
Large Language Model, Smart Grid, Threat Modeling, Prompt Injection
I Introduction
Artificial Intelligence (AI) technologies have been actively researched for different applications in smart grids. In recent years, various AI techniques ranging from classical Support Vector Machines (SVM) [1] to Artificial Neural Networks (ANN) [2] have been proposed for diverse applications within smart grids, including anomaly detection [2], load forecasting [3], energy theft detection [4], and so on.
Recently, a significant breakthrough in AI has been the development of Large Language Models (LLMs). LLMs have a large number of model parameters and possess the capability to process general natural language [5]. Due to their exceptional performance across various natural language processing (NLP) tasks, LLMs have facilitated numerous applications across different industries [6, 7]. Given the successful performance of AI techniques in numerous smart grid applications, it is natural to explore the potential of integrating LLMs into the smart grid [8, 9, 10, 11].
Figure 1: Illustration of cyberattacks to LLMs used in smart grid. The figure shows two cases: (1) an outsider attacker injects false information into LLMs and (2) an internal attacker obtains unauthorized information from LLMs.
Despite the outstanding performance of AI models, they can also be vulnerable and expose systems to new types of cyberattacks. Previous research has shown that popular AI models, such as deep neural networks, are susceptible to various attacks, including adversarial [12] and backdoor attacks [13]. These attacks have proven to be effective against AI applications in smart grids as well [14, 15, 16, 17]. Therefore, conducting thorough research on potential risks is crucial before deploying LLMs in smart grid applications.
The applications of LLMs are rapidly evolving, and research into their risks, especially in potential smart grid applications, is still nascent. To the best of our knowledge, [10] represents the sole research focusing on the security of LLMs in smart grid contexts; it outlines potential threat types but lacks a comprehensive threat model and validation. To address this gap, we systematically study the potential threats posed by LLMs in smart grid applications, develop generalized threat models, and validate these threats through simulated LLM attacks. Our contributions are summarized as follows:
  • We analyze how the threats posed by LLMs differ from those associated with previous AI models (e.g. SVM) and identify the scope of risks that LLMs bring to the smart grid.
  • Considering attack motivations, we propose two general threat models for potential attacks on LLMs deployed in smart grids: (1) bad data injection attacks by external adversaries, and (2) domain knowledge extraction attacks by internal actors.
  • As a proof of concept, we validate these two threat types using popular LLMs: OpenAI GPT-3.5 [18], OpenAI GPT-4 [19], and Meta LLaMA-3 [20]. Through thorough simulations, we demonstrate that attackers can readily inject malicious data into, and extract confidential information from, these models.
  • We have open-sourced the complete data, code, and evaluation results used in this paper to encourage further research on this topic [21].
The rest of the paper is organized as follows. The related work is presented in Section II. Section III presents the threat analysis and modeling of LLM applications smart grid. We presented our validation result in Section IV. Finally, Section V discusses the future works and concludes the paper.
II Related Work
II-A Machine Learning and NLP in Power Grids
Machine learning techniques have been extensively studied and applied in various smart grid applications, including load forecasting [22], anomaly detection [23], energy theft detection [4], and cyberattack detection [2, 24, 25, 26].
In recent years, natural language processing (NLP) applications have also appeared in smart grids. In 2016, Sun et al. utilized social network information to detect power outages in specific areas [27]. They employed Bayesian models to analyze Twitter messages, achieving accurate detection results. Similarly, Wang et al. [23] collected public news and utilized a pre-trained language model to detect incidents at renewable energy facilities.
In addition, research into the application of LLMs to power systems has begun. Dong et al. [11] explored potential use cases of LLMs within power systems, demonstrating the capabilities of fine-tuned LLMs in pattern recognition tasks such as equipment damage detection and document analysis. Similarly, Huang et al. [9] examined the use of LLMs in specific power system operations, including optimal power flow (OPF) and electric vehicle (EV) scheduling. These studies highlight the potential utility of LLMs in enhancing the efficiency and reliability of smart grids.
Further research on the application of machine learning in power grids can be found in [28].
II-B Machine Learning Security in Smart Grid
While ML has achieved outstanding results in many smart grid tasks, its security remains a significant concern.
Well-performing ML models have been shown to be vulnerable to adversarial attacks, a phenomenon originally demonstrated in the field of computer vision [12]. This type of attack has also proven effective against ML-based smart grid applications. In 2018, Chen et al. first explored adversarial examples with both categorical and sequential applications in power systems [14]. Later, Li et al. [17] demonstrated that attackers could craft adversarial data capable of bypassing ML-based cyberattack detection while adhering to the physical constraints of the power grid. [29] showed that adversarial attackers could also circumvent ML-based energy theft detection models while reporting extremely low power consumption data.
At the time of this research, the study by Ruan et al. [10] is the only work investigating the potential risks of applying LLMs in power systems. This work identifies critical risks such as privacy invasion, deteriorated performance, and semantic divergence.
III Threat Analysis and Modeling
III-A Background: Prompt and LLM
An LLM is a pre-trained ML model capable of processing natural language tasks. To achieve different task goals, LLMs are typically provided with a natural language prompt that specifies the context of the task. As shown in Fig. 2, an LLM application usually starts with the developer providing a prompt to the LLM, which defines how the LLM should process user inputs. The LLM then responds to the user’s input based on the information and requirements specified in the prompt.
Figure 2: LLM Application Workflow
As presented in [11, 9], when applying LLMs to the smart grid, the prompt provided to the LLM is expected to include domain knowledge and task requirements. Without loss of generality, the LLM query can be represented as:
[…]
III-B Threat Analysis: Scope
While research has identified various vulnerabilities in LLMs and their susceptibility to different types of attacks [30], such as jailbreak [31], prompt injection [32], and privacy violations [33], not all vulnerabilities are relevant in the smart grid context. To investigate the potential threats LLMs may pose to the smart grid, it is crucial to emphasize the properties of LLMs that differ from previous AI applications and focus on their practical usage in smart grid scenarios. Therefore, the scope of concern should exclude:
  • General LLM Inherent Limitations: Attacks such as jailbreaks, which attempt to circumvent LLM restrictions to output sensitive content like personal information, represent general issues for all LLM applications and are not unique to smart grids.
  • Privacy Concerns from LLM Service Providers: Products like ChatGPT may utilize user input to train underlying LLMs, potentially causing data leaks [34]. In practice, LLMs deployed for smart grid applications should operate within private or regulated cloud environments, and input data should be protected by the privacy policies of both smart grid operators and LLM service providers.
  • Unauthorized Access to Public LLMs: This is a broader security issue for smart grid network management. Such issues should be addressed through general network control measures, such as firewalls [35], rather than being seen as unique challenges posed by LLMs.
III-C Threat Models
Although not explicitly mentioned, most AI methods proposed for smart grid applications in previous literature assume that the AI model is trained by trusted personnel, such as internal engineers, and that all data used for training comes from reliable sources like smart meter data [22, 23, 4, 2]. However, this assumption no longer holds when applying LLMs. Due to the extremely high data requirements and costs associated with LLM training, pre-trained models are typically used. These pre-trained LLMs, whether open-source (e.g., Meta LLaMA-3) or proprietary (e.g., OpenAI GPT-4), have pre-trained parameters that may lead to unexpected behavior in response to well-crafted inputs from attackers.
Based on our analysis of existing research on LLM vulnerabilities, we categorize the main risks of LLMs for smart grid applications into bad data injection and domain knowledge extraction, as demonstrated in Figure 1. Accordingly, we propose two different threat models.
III-C1 Bad Data Injection
Bad data injection primarily concerns scenarios where LLMs receive input from public resources and are used to extract and process information from that input. This scenario is relevant to many practical applications within smart grids, such as processing user feedback reports from diverse sources.
We specifically consider cases where attackers attempt to inject bad data through public access points to the LLMs by manipulating legitimate inputs. The objective of such attacks is to deceive the LLMs into producing incorrect results that could mislead smart grid operators. For instance, an attacker could deceive the LLMs into misclassifying a report about a blackout as a thank-you letter by injecting well-crafted text into it. The general representation of bad data injection is
[…]
Here, we present the threat model for bad data injection targeting LLMs:
  • Public Access: LLMs receive inputs from a public access point, making them accessible to a wide user base.
  • Attacker Capabilities: The attacker is assumed to be able to modify legitimate input sent to the LLMs via this public interface. However, attackers cannot receive the direct output from the LLMs.
  • Attacker Knowledge: The attacker possesses knowledge about the high-level purpose of the LLMs and their potential output requirements.
Given that LLMs are utilized for processing information from publicly available sources, the existence of a public access point is inherent. Injecting bad data into legitimate inputs can be implemented in many ways, such as man-in-the-middle attacks. The high-level functions of the LLMs are sometimes disclosed to the public, like news and articles [36], which further facilitates attacker understanding and planning. Such information might be obtained from publicly accessible documents like a power utility portal’s user agreement and privacy policy, or through simple inference.
III-C2 Domain Knowledge Extraction
Domain knowledge extraction addresses scenarios involving unauthorized access to the domain knowledge that serves as the prompt for LLM applications. In many practical use cases, LLMs are provided with prompts containing raw data information but are designed to provide only specific responses to users’ queries. For example, the domain knowledge may include smart grid configuration information, where the LLM’s task is to answer specific questions based on this information without disclosing the raw configuration details. Domain knowledge extraction occurs when an attacker attempts to deceive the LLM into exposing domain knowledge that should not be revealed. This can be represented as:
[…]
Here, we present the threat model for domain knowledge extraction:
  • Attacker Capabilities: An insider attacker can interact with the LLM, including sending inputs and receiving outputs.
  • Attacker Knowledge: The attacker has high-level domain knowledge, such as understanding the functions of the LLM.
We focus on insider attackers in the context of domain knowledge extraction, as best practices recommend avoiding the provision of sensitive information to publicly accessible LLMs. Thus, insider attackers, who have legitimate access, can interact with the LLM by sending inputs and receiving outputs. By analyzing these outputs, the attacker can derive detailed information about the domain knowledge, potentially extracting sensitive information not intended for disclosure.
IV Validation
In this section, we validate the two types of attacks—bad data injection and domain knowledge extraction—by simulating scenarios involving LLM applications in a smart grid context. We selected the proprietary models OpenAI GPT-3.5 and GPT-4, as well as the open-source model Meta LLaMA-3, for our study due to their popularity and extensive coverage in related research [11, 9, 37]. Our validation results demonstrate that, under the practical threat models proposed in this paper, attackers can successfully execute both types of attacks on potential smart grid LLM applications.
The evaluations of the OpenAI models were implemented through the official OpenAI API [38]. For Meta LLaMA-3, we deployed the LLaMA-3.1-8B-Instruct [39] model on Amazon Web Services (AWS) using Hugging Face Inference Endpoints. Due to page limitations, only essential results are presented here; the implementation details, such as the complete system prompts and the datasets used in this paper, can be found in [21].
IV-A Bad Data Injection
IV-A1 Application Introduction
Figure 3: Message Samples
To demonstrate the susceptibility of LLMs to bad data injection attacks, we study the use case presented in [23] where an AI model was trained and used to perform a binary classification task to determine whether a message from public sources is related to an incident of renewable energy facilities. In practice, utilizing such an application can facilitate the rapid identification of potential incidents, thereby mitigating potential damage and enhancing response efficiency.
The dataset in [23] comprises messages collected from public news sources, focusing on energy events. Originally, the messages in the raw dataset were tokenized and not specific to renewable energy facilities. We revert the tokens to texts and manually curated the records to specifically include those related to renewable energy facilities. The refined dataset contains 423 labeled messages. Examples from the dataset are presented in Fig. 3, with ‘Yes’ representing related and ‘No’ not related.
IV-A2 Attack Simulation
Our attack simulation is structured into two distinct phases:
TABLE I: LLMs Performance with Normal Inputs
Model Accuracy Precision Recall F1
OpenAI GPT-3.5 89.1% 87.8% 89.6% 88.7%
OpenAI GPT-4 93.8% 94% 93.1% 93.5%
Meta Llama-3 88.9% 85.4% 92.6% 88.8%
REI Classifier [23] 88.5% 89.0% 93.8% 91.4%
  • Phase 1 - Normal Inputs: We reproduce the binary classification task using the LLMs. We empirically drafted a system prompt to identify relevant messages, and the same prompt was used for all three models. A sample of the system prompt can be found in Fig. 4. In this phase, we queried all three LLMs with normal messages, and the final input to the LLMs can be represented as
    $\text{Final Input}=\text{System Prompt}||\text{Normal Input}$Final Input = System Prompt | | Normal Input
    We summarize the performance of the LLMs on the classification task of normal inputs in Table I. By comparing the four metrics, we observe that all the LLMs achieve similar or better performance compared with the classifier proposed in [23]. We note that the effort required for the LLMs to achieve this performance is only to craft a simple system prompt, which is minimal effort compared with previous methods that include collecting and processing data, designing the model, and training. This demonstrates the efficiency of LLMs and their vast potential for application in smart grids.
  • Phase 2 - Injection Inputs: In this phase, we simulated scenarios where the normal inputs were modified by the attacker through the injection of crafted text. Considering the different motivations of the attackers, they may choose to perform a false negative attack (converting all classification results to ‘No’), a false positive attack (converting all classifications to ‘Yes’), or a reverse attack (reversing the classification results). The corresponding injection texts are shown in Fig. 4. We simulated these three types of attack by appending the injection text after the normal input. The final input to the LLMs can be represented as
[…]
    Similar to Phase 1, we queried the LLMs with the same messages but interspersed them with the three types of bad data sequentially to observe the effects on LLM responses. The performances of the LLMs under bad data injection attacks are shown in Table. II. By comparing the performances of the LLMs on normal inputs and injection inputs, we observe a significant drop in associated metrics for all three LLMs under all three different attack scenarios. These evaluation results demonstrate that well-designed injection texts can deceive the LLMs into misclassifying the inputs, highlighting the vulnerability of LLMs to bad data injection attacks in potential smart grid applications.
Figure 4: Bad Data Injection Prompts Presentation
TABLE II: LLMs Performance under Bad Data Injection
Model Metric Inject-FP Inject-FN Inject-Rev
OpenAI GPT-3.5 Accuracy 47.7% 52.2% 33.5%
Precision 47.7% 0% 34%
Recall 100% 0% 41.5%
F1 64.6% 0% 37.4%
OpenAI GPT-4 Accuracy 48.9% 52.2% 43.4%
Precision 48.3% 0% 45.4%
Recall 100% 0% 91.1%
F1 65.1% 0% 60.1%
Meta LlaMA-3 Accuracy 47.8% 52.2% 33.8%
Precision 47.8% 0% 38.8%
Recall 100% 0% 66.8%
F1 64.6% 0% 49.1%
IV-B Domain Knowledge Extraction
IV-B1 Application Introduction
To demonstrate the vulnerability of LLMs to domain knowledge extraction attacks from insider attackers, we consider the use case where LLMs are employed by a utility company acting as a virtual assistant for internal employees. Specifically, the LLMs are prompted with raw smart grid data, which we consider as domain knowledge, but are intended to respond to employees with only aggregated information about the data. Under normal circumstances, the LLMs would reject requests to output detailed data.
In this paper, we utilize the advanced metering infrastructure (AMI) data provided by the U.S. Energy Information Administration (EIA) [40] to serve as the domain knowledge. This AMI data originates from the EIA-861M form, a monthly report that gathers data on electricity sales and revenue from a statistically selected sample of electric utilities across the United States. For simplicity, we select only six columns of the dataset: Month, Utility_Number, Name, State, AMR_Total, AMI_Total. The sample system prompt (domain knowledge) provided to LLMs is shown in Fig. 5. To evaluate the generalization of the attack, we provide different data from EIA in the system prompts for each query we conduct.
Figure 5: System Prompt (Domain Knowledge) for LLM-based Virtual Assistant
IV-B2 Attack Simulation
Our attack simulation is structured into two distinct phases:
  • Phase 1 Normal Query: We directly challenge the LLMs by requesting them to output all the data they possess. This phase involves 100 tests for all three LLMs, each employing different groups of data (domain knowledge) to assess the robustness of the LLMs’ data confidentiality mechanisms under straightforward query conditions.
    In our evaluation, all three models successfully adhered to their configuration by not providing detailed data across all 100 tests. This demonstrates their effectiveness in safeguarding sensitive information under straightforward querying conditions. A sample of normal query input and LLM response is presented in Fig. 6
  • Phase 2 Domain Knowledge Extraction: In the attack phase, we employ well-crafted queries designed to subtly coax the LLMs into revealing more information than they are configured to disclose. This phase also consists of 100 tests using the same groupings of data as in Phase 1 for all three LLMs, to measure the LLMs’ resilience against more sophisticated extraction techniques.
    When faced with well-crafted, nuanced inputs designed to elicit more information than they are configured to disclose, all three models failed to maintain their data prevention protocols in all tests. In these cases, all models disclosed all the provided data, highlighting a vulnerability in their ability to handle sophisticated domain knowledge extraction attacks. The sampling output of the LLMs is shown in Fig. 6.
    Our evaluation results indicate that attackers can successfully extract the domain knowledge from LLM applications in smart grid environments under the threat models we proposed.
Figure 6: Domain Knowledge Extraction Input and Output
V Discussion and Conclusion
Large language models with pre-trained parameters pose significant challenges in controlling and understanding their generated outputs. Additionally, the domain knowledge utilized in LLM tasks is valuable and necessitates enhanced protection. Therefore, it is crucial to evaluate potential risks before deploying LLMs in smart grid applications.
In this paper, we delineated the scope of LLM security for smart grid applications and identified two major types of attacks concerning data security—bad data injection—and data privacy—domain knowledge extraction. Our validation demonstrates that attackers can successfully execute both types of attacks using well-crafted inputs. Given the rapid development of LLMs, new attack vectors may emerge. Monitoring these potential risks, such as interactivity injection, should be a focus of future research.